I have an isoscles triangle with a base of 4ft and a height of 5 ft. How do I find the length of the other two sides?
Thanks

Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

use a^2+b^2=c^2 with half the base
2^2+5^2=c^2
4+25=29
sqrt =5.4 ft each side

Answer by edjones(8007)

you have a 2 right triangles. The hypotenuse of either one is the side of the isosceles triangle.
b^2+a^2=c^2
2^2+5^2
=4+25
=29
c=sqrt(29)' length of one of the other sides of the isosceles triangle.
.
Ed